Gr ce Period: 3 days left | Register N PRINTER VERSIONBACK NEXT Problem 6.17 [ Your answer is incorrect. Try again. William Brown borrowed $15,550 from a bank for three years. If the quoted rate (APR) is 7.30 percent, and the compounding is daily, what isthe effective annual interest rate (EAR)? (Round answer to 2 decimal places e.g 15.25%. Use 365 days for calculation.) Effective Annual interest rate 23.54% Question Attempts: 1 of 2 used SAVE FOR LATER SUBMIT ANSWER wa
Answer #1

EAR = (1+quoted rate/n)^n - 1

n= compounding period

=(1+7.3%/365)^365 - 1

=(1+0.02%)^365-1

=(1.0002)^365 - 1

=1.075722-1

=0.075722

ie 7.57%